OpenStudy (anonymous):

One more square root Q... I think... How would I do: square root of 162/y^6

OpenStudy (anonymous):

do you know what \(\sqrt{y^6}\) is ? by now it should not be hard

OpenStudy (anonymous):

as for \(\sqrt{162}\) you should think of \(162\) as \(81\times 2\) and then finding \(\sqrt{162}\) should also be easy let me know if you still need help

OpenStudy (anonymous):

I'm so confused b/c there isn't a number times itself that equals 6, and I don't know how I would set it up... I'm sooo confused. I know I should know this, but it's just not sticking with me. Thanks for your patience with my stupidity hahaha! <3

OpenStudy (anonymous):

the 6 is in the exponent, not in the base you are not looking for the square root of 6, you are looking for the square root of \(y^6\) what is half of 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Haha :)

OpenStudy (anonymous):

right and so what is \(\sqrt{y^6}\) ?

OpenStudy (anonymous):

3..???? So y^3 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sooo I am just looking for half of the exponent then? And if it were the cubed root, then it would be...?

OpenStudy (anonymous):

how about \(\sqrt{y^{22}}\) ?

OpenStudy (anonymous):

if it was the cubed root, it would be ... one third of the exponent

OpenStudy (anonymous):

Hmmm... that is where I am confused. If it were y^22, then would it be y^11 with a 2 somewhere??? haha I am so confused :(

OpenStudy (anonymous):

if it was \(\sqrt{y^{22}}\) then you get \(y^{11}\) that is all

OpenStudy (anonymous):

I think it is coming to me! Thank you!! So if it were the cubed root of y^22...?

OpenStudy (anonymous):

if it was \(\sqrt{y^{23}}\) then you would get \(y^{11}\sqrt{y}\) because 2 goes in to 23 11 times with a remainder of 1 and it it was \(\sqrt[3]{y^{22}}\) then it would be \(y^7\sqrt[3]{y}\) becuase 3 goes in to 22 7 times with a remainder of 1

OpenStudy (anonymous):

Brilliant! Thank you thank you thank you! Life saver.

OpenStudy (anonymous):

it is not too hard when you get the hang of it. try some and see good luck on your exam

OpenStudy (anonymous):

I had a hang of it a week ago, but then I had a huge break and I always forget. Thank you - I will need all the luck I can get! Soooo worried haha.